# Knowledge Graph Large Language Model (KG-LLM) for Link Prediction

### v5

@@ 4.2 Multi-hop Link Prediction without In-Context Learning
This section delves into the comparative analysis of the traditional approaches, ablation framework and KG-LLM framework in the context of non-In-Context Learning (non-ICL) testing scenarios. Table 2 illustrates the baseline performance compared to the results achieved by our innovative approach. Traditional approaches are shown in the top section of the table, ablation framework is in the middle section, and KG-LLM framework is in the bottom section.
Answer to Q1: The comparative results demonstrate that across all models, the implementation of the KG-LLM framework surpasses the traditional approaches and ablation framework across both evaluated datasets. This improved performance can be attributed to the KG-LLM framework’s utilization of knowledge prompts. These prompts enable LLMs to take advantage of the intricate network of relationships between entities and their interconnections within the KG. Furthermore, the process of Instruction Fine-Tuning (IFT) precisely tailors the models, honing their focus on the limited options. The evidence presented here underscores the efficacy of our KG-LLM framework, enriched with CoT and IFT, indicating its potential to significantly advance the domain of multi-hop link prediction tasks in real-world applications.
We can observe that the performance of traditional models is not ideal. This performance is likely attributable to the complexities associated with multi-hop link prediction. Unlike the straightforward task of predicting a direct link between two nodes, traditional methods now must consider all intermediate nodes to arrive at a conclusion. This additional complexity consequently diminishes their performance. It should be noted that the performance metrics of the LlaMa2 and Gemma models within the ablation framework exhibit good performance. This observation suggests that even absent the enhancements introduced by the KG-LLM framework, models can still perform multi-hop link predictions. Nevertheless, the pivotal challenge is overcoming complex multi-hop KG problems. The KG-LLM framework emerges as a decisive factor in this area.
@@ 4.3 Multi-hop Link Prediction with In-Context Learning
To evaluate the influence of In-Context Learning (ICL) on models subjected to both the ablation and the KG-LLM frameworks, exclude the traditional approach as it lacks ICL capability. We conducted a comparative analysis using the same LLMs and testing inputs as in the previous section. The key distinction in this evaluation was the incorporation of an ICL example prior to each original testing input. This addition was meticulously designed to ensure task fairness and unbiased assessment; the ICL example shown in the appendix, derived from the training dataset, were restricted to a complexity level of three. This constraint aimed to prevent providing additional knowledge through the ICL example while furnishing a contextually relevant example. By adhering to these principles, we ensured a robust and fair evaluation of model performance under ICL conditions.
Table 3 reveals a notable enhancement in the performance of models under the ablation framework, with LlaMa2 and Gemma models achieving an F1 and AUC score exceeding 80%. Remarkably, the adoption of ICL within the KG-LLM framework resulted in a significant performance uplift; notably, the Gemma model achieved a staggering 98% F1 score on the first dataset, while LlaMa2 recorded a 96% F1 score on the second dataset.
An interesting observation emerged regarding the Flan-T5 model under the ablation framework on the second dataset, where performance did not improve or even slightly declined post-ICL integration. This phenomenon could be attributed to the increased length and complexity of the testing prompts induced by the added ICL context. While the inclusion of an ICL example generally aids model understanding, in certain cases, it might inadvertently be perceived as noise or irrelevant, potentially affecting the performance of the ablation framework Flan-T5.
Answer to Q2: The experimental results suggest that the deployment of ICL does not uniformly improve performance across all models within the ablation framework. However, for models refined through the KG-LLM framework, the integration of ICL consistently facilitates performance improvements. This finding underscores the nuanced role of ICL in model optimization, highlighting how effective it is in improving model performance within KG-LLM framework.
Table 4: Multi-hop Relation Prediction w/o In-Context Learning
Datasets WN18RR NELL-995
Metric Accuracy Accuracy
Flan-T5-Large (ablation) 0.14 0.06
LlaMa2-7B (ablation) 0.23 0.17
Gemma-7B (ablation) 0.22 0.19
Flan-T5-Large (KG-LLM) 0.29 0.21
LlaMa2-7B (KG-LLM) 0.33 0.25
Gemma-7B (KG-LLM) 0.36 0.27
Table 5: Multi-hop Relation Prediction with In-Context Learning
Datasets WN18RR NELL-995
Metric Accuracy Accuracy
Flan-T5-Large (ablation) 0.35 0.24
LlaMa2-7B (ablation) 0.56 0.48
Gemma-7B (ablation) 0.58 0.52
Flan-T5-Large (KG-LLM) 0.55 0.53
LlaMa2-7B (KG-LLM) 0.74 0.70
Gemma-7B (KG-LLM) 0.76 0.71
@@ 4.4 Multi-hop Relation Prediction without In-Context Learning
In this analysis, we explore the models’ ability to predict relations not encountered during their training phase, exclude the traditional approach as it lacks generalization ability. We maintain consistency with the testing dataset used in the multi-hop link prediction task to ensure comparability and fairness. The difference lies in the modifications made to the instruction and prompt question presented to the model. Detailed specifications of these instructions are provided in the Appendix.
An intriguing observation emerged from the evaluation, shown in Table 4: all frameworks showcased limited performance in this task. Notably, the KG-LLM framework exhibited marginally superior performance relative to its ablation counterpart. Upon examining the predictive outcomes, we reveal a predominant inclination towards negative relation predictions over positive ones, where models generated a ‘no relation’ output for negative instances. This trend was consistent across models operating under both frameworks. The difference in performance between the two frameworks mainly came from the fact that the KG-LLM models were better at accurately identifying some positive relationships.
@@ 4.5 Multi-hop Relation Prediction with In-Context Learning
In addition to evaluating the performance of models on the relation prediction task without In-Context Learning (ICL), this study further explores the impact of incorporating ICL into the task. The ICL example is shown in the appendix. The comparative results, as illustrated in Table 5, reveal a significant improvement in the generalization abilities of the models both under the ablation framework and the KG-LLM framework, in contrast to the findings presented in Table 4. In particular, the LlaMa2 and Gemma models, when processed through the KG-LLM framework and augmented with ICL, achieved an accuracy exceeding 70% across both datasets.
Answer to Q4: The integration of ICL has markedly improved the models’ ability to engage and excel in unseen tasks, as evidenced by the substantial performance gains detailed in Table 5. The KG-LLM framework, in particular, exhibits a heightened ability to learn and utilize the contextual example provided by ICL, thus significantly improving the accuracy in predicting unseen relations.
@@ 5 Conclusions and Future Work
Our study introduces the Knowledge Graph Large Language Model Framework (KG-LLM Framework) as a promising solution for multi-hop generative prediction tasks in knowledge graph analysis. By leveraging techniques such as Chain of Thought (CoT) and Instruction Fine-tuning (IFT), models processed by KG-LLM framework have greatly enhanced the accuracy of predictions in KG-related tasks. Overall, the ablation framework models have displayed relatively lower performance in multi-hop link prediction tasks, and the introduction of in-context learning (ICL) has further hindered its performance. However, the KG-LLM framework models have demonstrated strong performance in multi-hop link prediction tasks, both with and without the utilization of ICL. Additionally, it exhibits a moderate level of ability when aided by ICL in relation prediction tasks that are not seen during model pre-training. These findings highlight the importance of incorporating CoT and IFT methodologies in enhancing the performance of models, while also emphasizing the need for further exploration and improvement in relation prediction tasks and the ability of handling unseen tasks in general.
In our future work, we will consider accessing the model’s reasoning process in the evaluation stage. We are committed to further enhancing KG-related prediction tasks. One key aspect we will focus on is refining the instruction process by limiting the option size, which will help the model better understand and interpret the provided instructions. Additionally, we plan to explore the utilization of prompt quality filters to effectively filter out noisy options, improving the overall accuracy and reliability of the model’s predictions. Furthermore, we will dedicate efforts to advancing our model’s ability to handle tasks that are unseen during pre-training, aiming to enhance its capacity to generalize and handle unseen prompts more effectively. Through these ongoing improvements, we aim to advance the capabilities of KG-LLM models and contribute to the progress of knowledge graph analysis.
Appendix
5.1 Limitations
One limitation of the KG-LLM framework is that we have not tested the model’s performance when the training input exceeds 512 tokens. We anticipate that tasks will become considerably more challenging when dealing with datasets containing numerous relations. Since these relations must be incorporated as options in the instructions, the sheer volume of options can become extensive and potentially mislead the model. To overcome this limitation, future research endeavors to investigate techniques for effectively reducing the size of options. This endeavor will further enhance the efficacy of our KG-LLM framework approach.
5.2 Instruction for Multi-hop Link Prediction
### Instruction:
Below is the detail of a knowledge graph path. Is node [first node id] connnected with node [last node id]?
Answer the question by reasoning step-by-step.
Choose from the given options:
1. Yes
2. No
5.3 Instruction for Multi-hop Relation Prediction
### Instruction:
Below is the detail of a knowledge graph path. What is the relationship between the node [first node id] and the node [last node id]?
Answer the question by reasoning step-by-step.
Choose from the given options:
1. [relation 1]
2. [relation 2]
3. [relation 3]
5.4 Dataset 1
5.4.1 ablation Framework One-shot In-Context Learning for Multi-hop Link Prediction
@@ 5.4 Dataset 1
node_47405 has relation_179 with node_46497. node_46497 has relation_180 with node_46501. Is node 47405 connected with node 46501?
@@
5.4.2 KG-LLM Framework One-shot In-Context Learning for Multi-hop Link Prediction
@@
node_47405 has relation_179 with node_46497. node_46497 has relation_180 with node_46501. Is node 47405 connected with node 46501?
@@
node_47405 has relation_179 with node_46497, means node_47405 music_artist_genre node_46497. node_46497 has relation_180 with node_46501, means node_46497 music_genres_such_as_music_genres node_46501. So node 47405 music_artist_genre node 46501. The answer is yes.
5.4.3 ablation Framework One-shot In-Context Learning for Multi-hop Relation Prediction
@@
The relationship between the first node and the last node is relation_179.
5.4.4 KG-LLM Framework One-shot In-Context Learning for Multi-hop Relation Prediction
@@
node_47405 has relation_179 with node_46497, means node_47405 music_artist_genre node_46497. node_46497 has relation_180 with node_46501, means node_46497 music_genres_such_as_music_genres node_46501. So node 47405 music_artist_genre node 46501. The relationship between the first node and the last node is music_artist_genre.
5.5 Dataset 2
5.5.1 ablation Framework One-shot In-Context Learning for Multi-hop Link Prediction
### Context:
node_11809 has relation_9 with node_12218. node_12218 has relation_1 with node_9431. Is node 11809 connnected with node 9431?
Answer:
The answer is yes.
5.5.2 KG-LLM Framework One-shot In-Context Learning for Multi-hop Link Prediction
### Context:
node_11809 has relation_9 with node_12218. node_12218 has relation_1 with node_9431. Is node 11809 connnected with node 9431?
Answer:
node_11809 has relation_9 with node_12218, means node_11809 _verb_group node_12218. node_12218 has relation_1 with node_9431, means node_12218 _derivationally_related_form node_9431. So node 11809 _derivationally_related_form node 9431. The answer is yes.
5.5.3 ablation Framework One-shot In-Context Learning for Multi-hop Relation Prediction
### Context:
node_11809 has relation_9 with node_12218. node_12218 has relation_1 with node_9431. What is the relationship between node_11809 and node_9431?
Answer:
The relationship between the first node and the last node is relation_1.
5.5.4 KG-LLM Framework One-shot In-Context Learning for Multi-hop Relation Prediction
### Context:
node_11809 has relation_9 with node_12218. node_12218 has relation_1 with node_9431. What is the relationship between node_11809 and node_9431?
Answer:
node_11809 has relation_9 with node_12218, means node_11809 _verb_group node_12218. node_12218 has relation_1 with node_9431, means node_12218 _derivationally_related_form node_9431. So node 11809 _derivationally_related_form node 9431. The relationship between the first node and the last node is _derivationally_related_form.
5.6 Examples
Test Input Models’ Responses
Example of Multi-hop Link Prediction Without ICL
###Input:
node_60981 has relation_181 with node_9219. node_9219 has relation_146 with node_31785. node_31785 not connected with node_17868. node_17868 has relation_136 with node_17869. node_17869 not connected with node_23650. node_23650 has relation_11 with node_2407. Is node 60981 connected with node 2407?
Answer from ablation FLAN-T5-Large Model
The answer is no.
Example of Multi-hop Link Prediction With ICL
###Context:
node_47405 has relation_179 with node_46497. node_46497 has relation_180 with node_46501. Is node 47405 connected with node 46501?
Answer:
The answer is yes.
###Input:
node_20326 has relation_12 with node_20328. node_20328 not connected with node_8777. node_8777 has relation_71 with node_2906. node_2906 not connected with node_22546. node_22546 has relation_143 with node_10897. node_10897 has relation_2 with node_20. Is node 20326 connected with node 20?
Answer from ablation FLAN-T5-Large Model
The answer is no.
Example of Multi-hop Relation Prediction Without ICL
###Input:
node_57359 has relation_160 with node_57261. node_57261 has relation_161 with node_35831. node_35831 has relation_160 with node_35829. What is the relationship between node_57359 and node_35829 ?
Answer from ablation FLAN-T5-Large Model
The relationship between the first node and the last node is relation_160.
Example of Multi-hop Relation Prediction With ICL
###Context:
Answer:
The relationship between the first node and the last node is relation_179.
###Input:
node_58780 has relation_89 with node_5996. node_5996 has relation_1 with node_3. node_3 has relation_11 with node_17475. What is the relationship between node_58780 and node_17475 ?
Answer from ablation FLAN-T5-Large Model
The relationship between the first node and the last node is no.
Table 6: ablation FLAN-T5-Large Model Responses to Modified Inputs
Test Input Models’ Responses
Example of Multi-hop Link Prediction Without ICL
###Instruction:
Below is the detail of a knowledge graph path. Is node 57359 connnected with node 35829? Answer the question by reasoning step-by-step. Choose from the given options:
1. Yes
2. No
###Input:
node_57359 has relation_160 with node_57261. node_57261 has relation_161 with node_35831. node_35831 has relation_160 with node_35829.
Answer from KG-LLM Framework FLAN-T5-Large Model
node_57359 has relation_160 with node_57261, means node_57359 plantincludeplant node_57261. node_57261 has relation_161 with node_35831, means node_57261 plantgrowinginplant node_35831. node_35831 has relation_160 with node_35829, means node_35831 plantincludeplant node_35829. So node 57359 plantincludeplant node 35829. The answer is yes.
Example of Multi-hop Link Prediction With ICL
###Context:
… Is node 47405 connected with node 46501?
Answer:
… So node 47405 musicartistgenre node 46501. The answer is yes.
###Instruction:
Below is the detail of a knowledge graph path …
1. Yes
2. No
###Input:
node_57359 has relation_160 with node_57261. node_57261 has relation_161 with node_35831. node_35831 has relation_160 with node_35829.
Answer from KG-LLM Framework FLAN-T5-Large Model
node_57359 has relation_160 with node_57261, means node_57359 plantincludeplant node_57261. node_57261 has relation_161 with node_35831, means node_57261 plantgrowinginplant node_35831. node_35831 has relation_160 with node_35829, means node_35831 plantincludeplant node_35829. So node 57359 plantincludeplant node 35829. The answer is yes.
Example of Multi-hop Relation Prediction Without ICL
###Instruction:
Below is the detail of a knowledge graph path. What is the relationship between node_9595 and node_6033 ? Answer the question by reasoning step-by-step. Choose from the given options:
1. motherofperson
###Input:
node_9595 has relation_117 with node_10011. node_10011 not connected with node_7650. node_7650 has relation_91 with node_6033.
Answer from KG-LLM Framework FLAN-T5-Large Model
node_9595 has relation_117 with node_10011, means node_9595 producesproduct node_10011. node_10011 not connected with node_7650 means there is no relationship between node_10011 and node_7650. node_7650 has relation_91 with node_6033, means node_7650 athleteplayssport node_6033. So there is no connection between node 9595 and node 6033. The relationship between the first node and the last node is no.
Example of Multi-hop Relation Prediction With ICL
###Context:
… What is the relationship between node_47405 and node_46501 ?
Answer:
… The relationship between the first node and the last node is musicartistgenre.
###Instruction:
Below is the detail of a knowledge graph path …
1. motherofperson
###Input:
node_47478 has relation_10 with node_1912. node_1912 not connected with node_48232. node_48232 has relation_139 with node_46613.
Answer from KG-LLM Framework FLAN-T5-Large Model
node_47478 has relation_10 with node_1912, means node_47478 latitudelongitude node_1912. node_1912 not connected with node_48232 means there is no relationship between node_1912 and node_48232. node_48232 has relation_139 with node_46613, means node_48232 musicianinmusicartist node_46613. So there is no connection between node 47478 and node 46613. The relationship between the first node and the last node is no.
Table 7: KG-LLM Framework FLAN-T5-Large Model Responses to Modified Inputs
Test Input Models’ Responses
Example of Multi-hop Link Prediction Without ICL
###Input:
node_21118 has relation_2 with node_5. node_5 has relation_1 with node_332. node_332 has relation_73 with node_2405. node_2405 has relation_194 with node_31362. node_31362 has relation_153 with node_22206. node_22206 has relation_154 with node_27734. node_27734 has relation_2 with node_10. Is node 21118 connnected with node 10?
Answer from ablation LlaMa2-7B Model
The answer is yes.
Example of Multi-hop Link Prediction With ICL
###Context:
node_47405 has relation_179 with node_46497. node_46497 has relation_180 with node_46501. Is node 47405 connected with node 46501?
Answer:
The answer is yes.
###Instruction:
###Input:
node_21567 has relation_141 with node_21540. node_21540 has relation_141 with node_21378. node_21378 has relation_141 with node_21482. node_21482 has relation_141 with node_21391. node_21391 has relation_141 with node_21449. Is node 21567 connnected with node 21449?
Answer from ablation LlaMa2-7B Model
The answer is yes.
Example of Multi-hop Relation Prediction Without ICL
###Input:
node_58493 has relation_85 with node_2420. node_2420 has relation_7 with node_24721. node_24721 has relation_145 with node_24722. What is the relationship between node_58493 and node_24722 ?
Answer from ablation LlaMa2-7B Model
The relationship between the first node and the last node is no.
Example of Multi-hop Relation Prediction With ICL
###Context:
Answer:
The relationship between the first node and the last node is relation_179.
###Input:
node_64967 has relation_37 with node_20088. node_20088 has relation_30 with node_55250. node_55250 has relation_115 with node_11327. What is the relationship between node_64967 and node_11327 ?
Answer from Ablation LlaMa2-7B Model
The relationship between the first node and the last node is subpartoforganization.
Table 8: Ablation LlaMa2-7B Model Responses to Modified Inputs
Test Input Models’ Responses
Example of Multi-hop Link Prediction Without ICL
###Instruction:
Below is the detail of a knowledge graph path. Is node 21567 connnected with node 21449? Answer the question by reasoning step-by-step. Choose from the given options:
1. Yes
2. No
###Input:
node_21567 has relation_141 with node_21540. node_21540 has relation_141 with node_21378. node_21378 has relation_141 with node_21482. node_21482 has relation_141 with node_21391. node_21391 has relation_141 with node_21449.
Answer from KG-LLM Framework LlaMa2-7B Model
node_21567 has relation_141 with node_21540, means node_21567 chemicalistypeofchemical node_21540. node_21540 has relation_141 with node_21378, means node_21540 chemicalistypeofchemical node_21378. node_21378 has relation_141 with node_21482, means node_21378 chemicalistypeofchemical node_21482. node_21482 has relation_141 with node_21391, means node_21482 chemicalistypeofchemical node_21391. node_21391 has relation_141 with node_21449, means node_21391 chemicalistypeofchemical node_21449. So node 21567 chemicalistypeofchemical node 21449. The answer is yes.
Example of Multi-hop Link Prediction With ICL
###Context:
… Is node 47405 connected with node 46501?
Answer:
… The answer is yes.
###Instruction:
Below is the detail of a knowledge graph path….
1. Yes
2. No
###Input:
node_9595 has relation_117 with node_10011. node_10011 not connected with node_7650. node_7650 has relation_91 with node_6033.
Answer from KG-LLM Framework LlaMa2-7B Model
node_9595 has relation_117 with node_10011, means node_9595 producesproduct node_10011. node_10011 not connected with node_7650 means there is no relationship between node_10011 and node_7650. node_7650 has relation_91 with node_6033, means node_7650 athleteplayssport node_6033. So there is no connection between node 9595 and node 6033. The answer is no.
Example of Multi-hop Relation Prediction Without ICL
###Instruction:
Below is the detail of a knowledge graph path…
1. motherofperson
###Input:
node_20539 has relation_6 with node_10925. node_10925 has relation_120 with node_10922. node_10922 has relation_42 with node_10924.
Answer from KG-LLM Framework LlaMa2-7B Model
node_20539 has relation_6 with node_10925, means node_20539 synonymfor node_10925. node_10925 has relation_120 with node_10922, means node_10925 companyalsoknownas node_10922. node_10922 has relation_42 with node_10924, means node_10922 organizationterminatedperson node_10924. So node 20539 mutualproxyfor node 10924. The relationship between the first node and the last node is relation_7.
Example of Multi-hop Relation Prediction With ICL
###Context:
… What is the relationship between node_47405 and node_46501 ?
Answer:
… The relationship between the first node and the last node is musicartistgenre.
###Instruction:
Below is the detail of a knowledge graph path…
1. motherofperson
###Input:
node_40422 has relation_170 with node_59706. node_59706 has relation_170 with node_59850. node_59850 has relation_170 with node_59582.
Answer from KG-LLM Framework LlaMa2-7B Model
node_40422 has relation_170 with node_59706, means node_40422 professionistypeofprofession node_59706. node_59706 has relation_170 with node_59850, means node_59706 professionistypeofprofession node_59850. node_59850 has relation_170 with node_59582, means node_59850 professionistypeofprofession node_59582. So there is no connection between node 40422 and node 59582. The relationship between the first node and the last node is no.
Table 9: KG-LLM Framework LlaMa2-7B Model Responses to Modified Inputs
Test Input Models’ Responses
Example of Multi-hop Link Prediction Without ICL
###Input:
node_9595 has relation_117 with node_10011. node_10011 not connected with node_7650. node_7650 has relation_91 with node_6033. Is node 9595 connected with node 6033?
Answer from Ablation Gemma-7B Model
The answer is no.
Example of Multi-hop Link Prediction With ICL
###Context:
node_47405 has relation_179 with node_46497. node_46497 has relation_180 with node_46501. Is node 47405 connected with node 46501?
Answer:
The answer is yes.
###Input:
node_25714 has relation_1 with node_9667. node_9667 has relation_195 with node_2421. node_2421 has relation_73 with node_26004. node_26004 not connected with node_66961. node_66961 has relation_196 with node_24014. node_24014 has relation_1 with node_2156. node_2156 has relation_7 with node_25251. node_25251 not connected with node_4610. node_4610 has relation_56 with node_4806. node_4806 has relation_65 with node_2255. Is node 25714 connected with node 2255?
Answer from Ablation Gemma-7B Model
The answer is no.
Example of Multi-hop Relation Prediction Without ICL
###Input:
node_27123 has relation_97 with node_6069. node_6069 has relation_165 with node_6211. node_6211 has relation_165 with node_6163. What is the relationship between node_27123 and node_6163 ?
Answer from Ablation Gemma-7B Model
The relationship between the first node and the last node is no.
Example of Multi-hop Relation Prediction With ICL
###Context:
Answer:
The relationship between the first node and the last node is relation_179.
###Input:
node_49244 has relation_50 with node_1967. node_1967 has relation_50 with node_35902. What is the relationship between node_49244 and node_35902 ?
Answer from Ablation Gemma-7B Model
The relationship between the first node and the last node is no.
Table 10: Ablation Gemma-7B Model Responses to Modified Inputs
Test Input Models’ Responses
Example of Multi-hop Link Prediction Without ICL
###Instruction:
Below is the detail of a knowledge graph path. Is node 58493 connected with node 24722? Answer the question by reasoning step-by-step. Choose from the given options:
1. Yes
2. No
###Input:
node_58493 has relation_85 with node_2420. node_2420 has relation_7 with node_24721. node_24721 has relation_145 with node_24722.
Answer from KG-LLM Framework Gemma-7B Model
node_58493 has relation_85 with node_2420, means node_58493 politicianrepresentslocation node_2420. node_2420 has relation_7 with node_24721, means node_2420 mutualproxyfor node_24721. node_24721 has relation_145 with node_24722, means node_24721 cityliesonriver node_24722. So there is no connection between node 58493 and node 24722. The answer is no.
Example of Multi-hop Link Prediction With ICL
###Context:
… Is node 47405 connected with node 46501?
Answer:
… The answer is yes.
###Instruction:
Below is the detail of a knowledge graph path…
1. Yes
2. No
###Input:
node_58493 has relation_85 with node_2420. node_2420 has relation_7 with node_24721. node_24721 has relation_145 with node_24722.
Answer from KG-LLM Framework Gemma-7B Model
node_58493 has relation_85 with node_2420, means node_58493 politicianrepresentslocation node_2420. node_2420 has relation_7 with node_24721, means node_2420 mutualproxyfor node_24721. node_24721 has relation_145 with node_24722, means node_24721 cityliesonriver node_24722. So there is no connection between node 58493 and node 24722. The answer is no.
Example of Multi-hop Relation Prediction Without ICL
###Instruction:
Below is the detail of a knowledge graph path…
1. motherofperson
###Input:
node_1946 has relation_1 with node_3. node_3 has relation_7 with node_346. node_346 has relation_1 with node_2572.
Answer from KG-LLM Framework Gemma-7B Model
node_1946 has relation_1 with node_3, means node_1946 proxyfor node_3. node_3 has relation_7 with node_346, means node_3 mutualproxyfor node_346. node_346 has relation_1 with node_2572, means node_346 proxyfor node_2572. So node 2572 atlocation node 1946. The relationship between the first node and the last node is relation_11.
Example of Multi-hop Relation Prediction With ICL
###Context:
… What is the relationship between node_47405 and node_46501 ?
Answer:
… The relationship between the first node and the last node is musicartistgenre.
###Instruction:
Below is the detail of a knowledge graph path…
1. motherofperson
###Input:
node_42306 has relation_2 with node_21. node_21 has relation_1 with node_332.
Answer from KG-LLM Framework Gemma-7B Model
node_42306 has relation_2 with node_21, means node_42306 atdate node_21. node_21 has relation_1 with node_332, means node_21 proxyfor node_332. So node 332 locationlocatedwithinlocation node 42306. The relationship between the first node and the last node is locationlocatedwithinlocation.
Table 11: KG-LLM Framework Gemma-7B Model Responses to Modified Inputs

### v6

@@ 4.2 Multi-hop Link Prediction without In-Context Learning
This section analyzes the traditional approaches, ablation framework, and KG-LLM framework in the context of non-In-Context Learning (non-ICL) Link Prediction, as shown in Table 2. Traditional approaches are shown in the top section of the table, the ablation framework is in the middle section, and the KG-LLM framework is in the bottom section.
Answer to Q1: Our analysis reveals that the traditional approach’s GNN model, especially ConvRot, exhibited relatively good performance, particularly surpassing the ablation models on the WN18RR dataset. This GNN model performance can be attributed to its ability to effectively capture the structural information in graph data. However, the results demonstrate that across all models, the implementation of the KG-LLM framework surpasses the traditional approaches and ablation framework across all datasets. This improved performance can be attributed to the KG-LLM framework’s knowledge prompts. These prompts enable LLMs to take advantage of the relationships network between entities and their interconnections within the KG. Furthermore, these LLMs already possess basic common sense knowledge from pre-training. When all nodes and relations are converted to text, this inherent common sense enhances their understanding of the relations and nodes, thereby improving link prediction accuracy. Instruction fine-tuning (IFT) also contributed to this improvement by forcing models to focus on the limited options. The evidence presented here underscores the efficacy of our KG-LLM framework, enriched with CoT and IFT, indicating its potential to advance the domain of multi-hop link prediction tasks in real-world applications.
We also evaluate GNN, ablation, and KG-LLM framework models’ performance at each level of hop complexity in WN18RR and NELL-995 datasets. As shown in Figure 4, the performance of GNN and ablation models significantly declines as hop complexity increases. Upon closer examination, it is evident that as hop complexity grows, these models frequently respond with ’No’ for most questions, resulting in an F1 score close to 0 and an AUC score around 0.5. This performance is due to the increased complexities of multi-hop link prediction. Unlike the straightforward task of predicting a direct link between two nodes, models must consider all intermediate nodes to conclude, adding significant complexity and reducing their effectiveness. In contrast, the KG-LLM framework models effectively address this challenge, maintaining fair performance even at five-hops, except for the Flan-T5 model.
@@ 4.3 Multi-hop Link Prediction with In-Context Learning
In this section, we evaluate the influence of In-Context Learning (ICL) on models subjected to both ablation and KG-LLM frameworks, excluding the traditional approach as it lacks ICL capability. We experimented using the same LLMs and testing inputs as in the previous section. The key distinction in this evaluation was adding an ICL example at the beginning of each original testing input. The ICL example shown in Appendix A.1, derived from the training dataset, was restricted to the complexity of two-hops. This constraint aimed to prevent providing additional knowledge through the ICL example while furnishing a contextually relevant example.
Table 3 reveals a notable enhancement in the performance of models under the ablation framework, with LLaMa2 and Gemma models achieving an F1 and AUC score exceeding 80% in WN18RR and NELL-995 datasets. Remarkably, the adoption of ICL within the KG-LLM framework resulted in a significant performance uplift. Notably, the Gemma model achieved a staggering 98% F1 score on the first dataset, while LlaMa2 recorded a 96% F1 score on the second dataset.
An interesting observation is that ICL has shown unstable improvements in the Flan-T5 model. For some datasets within the ablation and KG-LLM frameworks, the performance slightly declined after implementing ICL. This phenomenon could be attributed to the increased length and complexity of the testing prompts. While the inclusion of an ICL example generally aids in model understanding, in certain cases, it might be perceived as noise, potentially affecting the Flan-T5’s performance.
Answer to Q2: The experimental results indicate that the deployment of ICL does not uniformly improve performance across all models. However, for the LLaMa2 and Gemma models, the integration of ICL consistently facilitates performance improvements.
(a) Ablation Framework
(b) KGLLM Framework
Figure 5: Multi-hop Relation Prediction Performance Comparison: The left graph shows model performance under the ablation framework, while the right graph shows model performance under the KGLLM framework. Blue bars represent testing without ICL, and red bars represent testing with ICL.
@@ 4.4 Multi-hop Relation Prediction without In-Context Learning
In this analysis, we explore models’ ability to perform unseen multi-hop relation prediction tasks on WN18RR and NELL-995 datasets, excluding the traditional approach as it lacks generalization ability. We used the same testing dataset in the multi-hop link prediction task to ensure comparability and fairness. As mentioned in section 3.2, the difference lies in the instruction and prompt question presented to the model.
Our findings are presented in Table 5. We discovered that both frameworks showcased limited performance in this task without ICL. Notably, the KG-LLM framework exhibited marginally superior performance. Upon reviewing the predictive results, we observed that the model continues to provide ‘yes’ and ‘no’ answers for most questions, similar to the multi-hop link prediction task. For some questions, it outputs random responses.
@@ 4.5 Multi-hop Relation Prediction with In-Context Learning
We further explore the impact of incorporating ICL into the multi-hop relation prediction task. The ICL example is shown in the Appendix A.1. The results of red bars (with ICL) in Table 5 reveal a significant improvement in the generalization abilities of the models under both ablation and KG-LLM frameworks, in contrast to the blue bars (without ICL). In particular, the LlaMa2 and Gemma models under the KG-LLM framework with ICL, achieved an accuracy exceeding 70% in the WN18RR datasets.
Answer to Q4: The integration of ICL has improved the models’ ability to excel in unseen tasks. The KG-LLM framework, in particular, exhibits the ability to learn and utilize the contextual example provided by ICL.
@@ 5 Conclusions and Future Work
Our study introduces the Knowledge Graph Large Language Model Framework (KG-LLM Framework) as a promising solution for multi-hop generative prediction tasks in knowledge graph analysis. By leveraging techniques such as Chain of Thought and Instruction Fine-tuning, models processed by KG-LLM framework have greatly enhanced the accuracy of predictions in KG-related tasks.
In our future work, we will consider accessing the model’s reasoning process in the evaluation stage. We are committed to further enhancing KG-related prediction tasks. One key aspect we will focus on is refining the instruction process by limiting the option size. Additionally, we plan to explore the utilization of prompt quality filters to effectively filter out noisy options, improving the overall accuracy and reliability of the model’s predictions. Through these ongoing improvements, we aim to advance the capabilities of KG-LLM models and contribute to the progress of knowledge graph analysis.
Appendix A Appendix
A.1 In-Context Learning Examples
A.1.1 Multi-hop Link Prediction ICL Example in Ablation Framework
@@
Node_47405 has relation_179 with Node_46497. Node_46497 has relation_180 with Node_46501. Is Node_47405 connected with Node_46501?
@@
A.1.2 Multi-hop Link Prediction ICL Example in KG-LLM Framework
@@
Node_47405 has relation_179 with Node_46497. Node_46497 has relation_180 with Node_46501. Is Node_47405 connected with Node_46501?
@@
Node_47405 has relation_179 with Node_46497, means Miles Davis music artist is associated with genre Bebop. Node_46497 has relation_180 with Node_46501, means Bebop genre is under the broader genre Jazz. So Miles Davis music artist is associated with genre Jazz.
The answer is yes.
A.1.3 Multi-hop Relation Prediction ICL Example in Ablation Framework
@@
The relationship between the Node_47405 and Node_46501 is relation_179.
A.1.4 Multi-hop Relation Prediction ICL Example in KG-LLM Framework
@@
Node_47405 has relation_179 with Node_46497, means Miles Davis music artist is associated with genre Bebop. Node_46497 has relation_180 with Node_46501, means Bebop genre is under the broader genre Jazz. So Miles Davis music artist is associated with genre Jazz.
The relationship between Miles Davis and Jazz is music_artist_genre.